OpenStudy (anonymous):

Solve using the square root method. x2 - 16 = 0 A.{-8, 8} B.{-4, 4} C.{8} D.{4}

OpenStudy (eyust707):

okay so first add 16 to both sides

OpenStudy (eyust707):

what do you get?

OpenStudy (campbell_st):

x^2 = 16 \[x = \sqrt{16}\] \[x = \pm ?\]

OpenStudy (anonymous):

@ campbell_st x=4

OpenStudy (eyust707):

correct!

OpenStudy (eyust707):

its impoortant to remember tho what x^2 really means.

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

so the answer would be d then thanks very much

OpenStudy (campbell_st):

and remember the are 2 solutions to this problem

OpenStudy (eyust707):

nope not d

OpenStudy (callisto):

Nah, don't forget x = ±2, not 2 itself alone!

OpenStudy (eyust707):

check it out

OpenStudy (eyust707):

x^2 = x * x = 16 "what times what equals 16?"

OpenStudy (callisto):

Sorry x = ±4, not 4 itself alone

OpenStudy (eyust707):

well 4 times 4 equals 16

OpenStudy (anonymous):

oh there are two so it would be b

OpenStudy (eyust707):

but what about -4 times -4? does -4 times -4 equal 16?

OpenStudy (eyust707):

correct!

OpenStudy (anonymous):

but 8 * 8 doesn't equal 16 either...

OpenStudy (eyust707):

nope 8 * 8 = 64

OpenStudy (anonymous):

yes i know that ^

OpenStudy (eyust707):

okay so we start with x^2 -16 = 0

OpenStudy (eyust707):

we are then allowed to add the same thing to both sides

OpenStudy (eyust707):

+16 is a smart choice

OpenStudy (eyust707):

so add 16 to the left side and add 16 to the right side

OpenStudy (eyust707):

then we end up with x^2 = 16

OpenStudy (eyust707):

well x^2 just means x * x

OpenStudy (eyust707):

so x*x =16

OpenStudy (eyust707):

or "what times what equals 16?"

OpenStudy (anonymous):

4 * 4 = 16

OpenStudy (eyust707):

-4 * -4 = ?

OpenStudy (anonymous):

16 because a negative times a negative equals a positive

OpenStudy (eyust707):

yes!

OpenStudy (eyust707):

so we could plug 4 in for x and get 16 but we could also plug in -4!

OpenStudy (anonymous):

ypu

OpenStudy (eyust707):

so there are two numbers that work!

OpenStudy (anonymous):

yup*

OpenStudy (eyust707):

4 and -4 both work